A mass of 5.5 kg of saturated liquid-vapor mixture of water is contained in a piston-cylinder at 101.325 kPa. Initially, 2.3 kg of the water is in the liquid phase and the rest is in the vapor phase. Heat is now transferred to the water, and the piston, which is resting on a set of stops, starts moving when the pressure inside reaches 300 kPa. Heat transfer continues until the total volume increases by 21 percent. Determine:
a) the initial temperature (oC),
b) the final temperature (oC),
c) the specific volume of water when the piston first starts moving (m3/kg), and
d) the work done during this process (kJ).

T1 (oC) Format : 94.62
T3 (oC) Format : 227.2
v2 (m3/kg) Format : 0.945
Wb (kJ) Format : 242.2
Answer #1

Pressure Voleme (P-V) of the system is as follows- 23 :582 + We have 1 given total mass = m=5.5kg mass of liquid= mg=2-3 kg sspecific volume of saturated vapor Ug = 1-6733 m3/kg specific internal energy of saturated vapor Ug = 2505.9848 k slka. Parto- - We have given pressure al starte 3 as P3= 3ookpa . So that final temperature T3 is given from steam table at P3 and ils T